# Computational Analysis of Character Development in Holocaust Testimonies

### v1

@@ Abstract
This work presents a computational approach to analyze character development along the narrative timeline. The analysis characterizes the inner and outer changes the protagonist undergoes within a narrative, and the interplay between them. We consider transcripts of Holocaust survivor testimonies as a test case, each telling the story of an individual in first-person terms. We focus on the survivor’s religious trajectory, examining the evolution of their disposition toward religious belief and practice along the testimony. Clustering the resulting trajectories in the dataset, we identify common sequences in the data. Our findings highlight multiple common structures of religiosity across the narratives: in terms of belief, most present a constant disposition, while for practice, most present an oscillating structure, serving as valuable material for historical and sociological research. This work demonstrates the potential of natural language processing techniques for analyzing character evolution through thematic trajectories in narratives.
@@ 1 Introduction
Characters in narratives are shaped by the continuous interplay of internal and external events they experience, each influencing their thoughts, feelings, and behaviors. As narratives progress, these events unfold along various dimensions. Some are common to essentially all narratives, such as changes in location, while others are specific to particular genres. Within the domain of Holocaust testimonies, as people experience unimaginable atrocities, the change in religious dispositions plays a central role Brenner (1980). Applying NLP technology for analyzing these testimonies has been advocated (Artstein et al., 2016; Wagner et al., 2022), offering the potential to extract valuable insights from the vast number of testimonies, instead of relying on small-scale, manual studies.
In this work we extract representations of this progression along the narrative timeline. This allows us to analyze religious trajectories across the entire corpus of testimonies, rather than focusing on individual cases. We leverage off-the-shelf LLMs for inferring the character arc of the protagonist, focusing on their development by exploring the progression of their religious trajectory throughout the narrative. We conduct a dual perspective analysis that captures the protagonist’s descriptions of their beliefs, and religious practices.
As people change and adapt, their observance of religion and their disposition towards religion constantly evolve. This is particularly relevant when discussing Holocaust testimonies, which follow survivors’ experiences in extreme conditions. We expect different patterns of change in religiosity throughout the testimonies, each telling the story of an individual and covering a wide range of times and places in their life.
We define a religious trajectory as a series of mentions of religiosity throughout a given storyline of an individual’s life course. We view each trajectory as a combination of the evolution of religious practices and beliefs, assuming that they each describe religiosity from a unique perspective.
@@ 2 Related Work
Recent NLP research has underscored the importance of narrative analysis in the study of human behavior and belief systems. While these studies often center around event and location schemas Wagner et al. (2023), character personalities, narrator intentions Piper et al. (2021); Zhu et al. (2023), or narrative structure Wagner et al. (2022), there remains a gap in tracking the evolution of characters, particularly their development. This gap is critical, as understanding character evolution can provide deeper insight into how narratives convey complex human experiences.
@@
This study builds upon these foundations by analyzing character development, focusing on the protagonist’s religious practices and their expressed religious belief systems.
@@
In the context of the Holocaust, Lassley (2015) examines religious transformation among Jewish survivors of the Holocaust, focusing on testimonies that describe a shift from being connected with religious traditions to abandoning belief in God completely. Their work discusses how survivor testimonial information can serve as valuable sources regarding faith after the Holocaust.
@@ 3 Task Definition
Given a segmented testimony $(x_{1},x_{2},\ldots,x_{n})$, our framework outputs a subsequence of segments representing its religious trajectory $(\langle b_{1},p_{1},t_{1}\rangle,\langle b_{2},p_{2},t_{2}\rangle,\ldots,\langle b_{k},p_{k},t_{k}\rangle)$, where $b_{j}$ are the labels for the belief trajectory, $p_{j}$ for practice, and $t_{j}\in(0,1)$ represents their position within the testimony.
@@ 4 Data and Annotation
We received 1000 Holocaust survivor testimony transcripts archived by the Shoah Foundation (SF). [^1] https://sfi.usc.edu/ The interviews were recorded on video in English and transcribed as time-stamped text. The significance of this dataset in the study and remembrance of the Holocaust cannot be overstated. As the number of surviving witnesses diminishes, there is an urgent need to find new approaches to engage with the extensive collection of testimonies.
Due to the lack of a structured format, our first step is to segment the data using natural divisions created by question-answer pairs followed by merging segments with fewer than 10 words and dividing those exceeding 100 words. The method aims to create segments that contain just enough information to determine its relevance to the religious trajectory. Our segmentation method yields a dataset, with the majority length of the segments being between 50-100 words and reasonable paragraph breaks, mostly keeping separate ideas apart from each other. However, some of the samples may contain multiple topics or conflicting religious signals. For example: “Saturday I would go, and I was encouraged to go to the synagogue. And Sunday, I was encouraged to go to church.”
@@ 4.1 Annotations
After segmenting the data, we randomly sample and annotate approximately 4000 segments. The trajectories are constructed from all religious content segments from the narrator’s perspective. The first annotation assignment is for capturing descriptions of Jewish religious practices and beliefs. This is designed as a binary classification task – the annotators are directed to classify all segments that describe any Jewish religious practices or beliefs of the protagonist or indicate their absence. We provide the complete guidelines in Appendix A.
@@
This complex annotation task involves knowledge of history and the Jewish religion. Many segments can be ambiguous, rely on context, or require an understanding of the context. Therefore, there may be a lack of consensus among the annotators as to the exact classification of a given segment. All these factors contribute to the complexity of the annotation task and require careful consideration and attention.
To enhance the accuracy of the annotation process and the quality of its results, we recruit multiple annotators, each tasked with annotating over 2500 segments, with an overlap of about 500, for measuring inter-annotator agreement (IAA). The overlap samples are selected randomly and the annotators did not know of the overlap. To evaluate the IAA, we use Krippendorff’s alpha, obtaining a score of 0.64, indicating a substantial level of agreement.
@@
The second annotation task is multi-label and consists of identifying the class of each sample and determining the speaker’s valence toward the specific class. Each segment belonging to the trajectory is categorized by its content – practice, belief, or both. The segments are further categorized by their valence toward the specific practice or belief they describe. Specifically, we map each sample to one or more of the classes or mark it as falsely recognized religious content.
We distinguish between religious practices and beliefs, as this is discussed in the Sociology of Religion. Chaves (2010) describes common incongruities among religious observers. He states: “Ideas and practices exist as bits and pieces that come and go as situations change, producing many inconsistencies… Religious ideas, values, and practices generally are not congruent.”
@@
The classification schema is as follows: a segment describing a practice can be annotated with one of the classes – Active, Inactive or Other. The first two classes identify descriptions of participating or violating religious practices while Other represents descriptions that do not meet the criteria of Active or Inactive, or match both of the classes simultaneously. For belief, the classification schema is similar, with the classes being Positive, Negative, and Other. An example of the annotation process is illustrated in Figure 2, and the annotation guidelines are detailed in Appendix B.
@@
To construct the dataset for this task, we first run the religious content classifier to select positive predictions. Next, we use GPT-4o [^2] https://platform.openai.com/docs/models/gpt-4o version gpt-4o-2024-05-13 to identify samples that belong to the minority classes, specifically the Belief classes and Other-Practice. The prompts used with GPT-4o include definitions of these minority classes. We then incorporate into the dataset all samples that GPT-4o classified as belonging to these minority classes. This approach helps balance the dataset by increasing the number of examples in the underrepresented classes, ensuring both positive and negative instances are well-represented. The resulting dataset is the one provided to the annotators.
@@
The Krippendorff’s alpha score for the second task is 0.44, indicating moderate agreement and suggesting that this classification task is challenging even for human annotators. The Belief dataset contains 141 Positive samples, 76 Negative, and 42 Other. For Practice predictions, the class distribution is 351 Active, 76 Inactive, and 44 Other. We divide the data into three splits with equal portions of each class.
@@ 5.1 Experimental setup
Extracting the religious trajectory of a given segmented testimony consists of two annotation tasks followed by prompting and training adapters for LLM classification. The main steps are: (1) filter religious content, (2) identify practices and beliefs, and (3) measure the valence of the protagonist’s expressed attitude toward religion.
@@
The annotation process yields approximately 700 positive samples with religious content and the rest negative. We fine-tune a RoBERTa-large Liu et al. (2019) classifier on this task, using a balanced dataset of approximately 1400 segments, divided into three splits: train, validation, and test, by a ratio of 0.8, 0.1 and 0.1, respectively. [^3] Hyperparameters: 3 epochs, batch size=8, learning rate=1e-5, seed=33 We evaluate the model on the test set, which yields accuracy and F1 scores of 95%. Additionally, cross-validation over 10 random splits produces an average accuracy of 92%. Given that the goal of this assignment is to identify segments that form a node in the trajectory – i.e., contain relevant practice and belief descriptions, these results demonstrate that the RoBERTa classifier is a reliable tool for this purpose.
For modeling the trajectories, we fine-tune LoRA adapters for Mistral-7B-instruct on data from the second annotation task and compare their performance with GPT-4o. Model selection was based on the average F1 score. For training, we use Mistralai’s mistral-finetune repository [^4] https://github.com/mistralai/mistral-finetune with their default hyperparameters and test on multiple seeds. [^5] Hyperparameters: Lora Rank=64, sequence length=16,384, batch size=1, learning rate=1e-4 number of micro-batches=1. Number of training steps for belief=100; for practice=400, weight decay=0.1, pct-start=0.05. Seed for belief=1; for practice=5 For the four practice classes, the average F1 scores are 0.51 for GPT-4o and 0.55 for Mistral-7B. For the parallel belief classes, the scores are 0.57 and 0.56, respectively. In a zero-shot setting, the average F1 scores for practice are 0.45 for GPT-4o and 0.33 for Mistral-7B. For the parallel belief classes, the scores are 0.56 and 0.23, respectively.
Both models perform comparably on our test data, leading us to select Mistral-7B for generating the trajectories. The prompts we run fit the format of Self-Consistency Wang et al. (2022) and are carefully designed with guidance from Anthropic’s prompt generator. [^6] https://console.anthropic.com/login Additionally, we tested GPT-4777version: gpt-4-turbo-2024-04-09 on a subset of the dataset, but it did not respond in the specified format. Considering the complicated annotation process and IAA, together with the models’ bias toward stereotypes in the context of religion Plaza-del Arco et al. (2024), these moderate F1 scores are unsurprising.
@@ 5.2 Evaluation
As we are not aware of any directly comparable reference for the trajectories. We therefore evaluate them against two reliable proxies, albeit not directly comparable ones: a topic-modeling-based and a thesaurus-based.
@@
Ifergan et al. (2024) conducted topic modeling using BERTopic, on the same set of testimony transcripts. Their results suggest multiple topics that align with our definitions of religious practices and beliefs. Using this prior knowledge, we extract the labeling by these topics as reference trajectories. Their segmentation method creates segments about three times larger than ours, leading to relatively sparse sequences. The topics we address, and their matching categories appear in Appendix E.
@@
The testimony transcripts we received from the SF were divided into segments according to their duration. Each segment was indexed with labels describing several terms from the SF thesaurus, containing $\sim$ 8000 unique labels.888sfi.usc.edu/content/keyword-thesaurus We leverage the terms related to religious practices and beliefs to extract reference trajectories. The full list of addressed terms can be found in Appendix F.
@@
Both of these references have several limitations. First, the segments in the reference trajectories are larger than ours, leading them to contain information irrelevant to the practice and belief labels. Second, the reference sequences are sparse compared to the predicted ones. For the topics, since our analysis for the topic model selects a single topic for each segment, rather than all the top relevant topics. For the thesaurus, despite the large number of terms, many segments lack the labeling of their matching terms. This may be a result of the exceedingly large label set ($\sim$ 8K labels), which may have resulted in recall issues in the SF annotation. In addition, the label set of the references is partial, meaning that the majority of the topics and terms that we address do not have a specific valence. Considering these points, the false positive rate of the predictions is unknown, making the references suitable only for measuring recall. An example illustrating the alignment of the references and predictions is shown in Figure 3.
@@
If $T$ is empty, return the number of segments in $R$, and if $R$ is empty, return 00.
We compare the sum of the distances to baseline trajectories. These baselines are artificial sequences designed by using our prior knowledge about the distribution (Figure 5) of practice and belief segments in the full dataset; each is the same length as the predicted sequence. The different baselines for a given class are defined in Table 1.
@@
Normal-original Sample from the Gaussian with the variance and mean values of the predictions.

### v5

@@ Abstract
This work presents a computational approach to analyze character development along the narrative timeline. The analysis characterizes changes in the protagonist’s views and behavior and the interplay between them. We consider transcripts of Holocaust survivor testimonies as a test case, each telling the story of an individual in first-person terms. We focus on the survivor’s religious trajectory, examining the evolution of their disposition toward religious belief and practice as it is reflected in the testimony. Clustering the resulting trajectories in the dataset, we identify common sequences in the data. Our findings highlight multiple common structures of religiosity across the narratives: in terms of belief, a constant disposition is common, while for practice, most present an oscillating structure, serving as valuable material for historical and sociological research. This work demonstrates the potential of natural language processing for analyzing character evolution through thematic trajectories in narratives.
@@ 1 Introduction
Characters in narratives are shaped by the continuous interplay of internal and external events they experience, each influencing their thoughts, feelings, and behaviors. As narratives progress, these events unfold along various dimensions. Some are common to essentially all narratives, such as changes in location, while others are specific to particular genres Eder (2010). Personal narratives, told in the first person, offer insights into the inner-life development of the protagonist McAdams (2001).
Within the domain of Holocaust testimonies, as people experience unimaginable atrocities, the change in religious dispositions plays a central role Brenner (1980). Schweid (1988) discusses the post-Holocaust crisis of theodicy in Jewish and Christian religious thought, necessitating substantial changes in religious content and norms. Specifically, the question of how religious faith evolves under extreme circumstances remains actively debated in Holocaust studies, with scholars like Berkovits (1973) stressing that the majority of survivors maintained their faith, and Leo et al. (2021) discussing the role of faith as a stabilizing force, arguing that most individuals do not change their religious beliefs following traumatic experiences. In contrast, Ben-Ezra et al. (2010) describe common processes of change in both directions. This ongoing argument makes Holocaust testimonies a suitable test case for religious evolution analysis, offering further insight by analyzing a large corpus of testimonies through a holistic computational approach that could extend to other dimensions of character development.
Holocaust survivor testimonies are indispensable for both scholarship and collective memory. Applying NLP to historical texts, and in particular to these testimonies, has been advocated (Artstein et al., 2016; Wagner et al., 2022), offering the potential to extract valuable insights from the vast number of testimonies, instead of relying on small-scale, manual studies (Presner et al., 2024; Garcia and Weilbach, 2023). Following this approach, we utilize NLP technology to examine religious trajectories across an entire corpus of testimonies, instead of focusing on individual cases. Leveraging off-the-shelf LLMs for inferring the character arc of the protagonist, we explore the progression of their religious trajectory throughout the narrative. We conduct a dual perspective analysis that captures the protagonist’s descriptions of their beliefs and religious practices.
@@
As people change and adapt, their disposition towards religion and level of observance constantly evolve. This is particularly relevant when addressing Holocaust testimonies, which follow survivors’ experiences in extreme conditions. We expect to reveal different patterns of change in religiosity throughout the testimonies, each telling the story of an individual and covering a wide range of periods and locations.
We define a religious trajectory as a series of mentions of religiosity throughout a given storyline of an individual’s life course. Separate trajectories are held for religious practices and beliefs, assuming they each describe religiosity from a unique perspective. This distinction is discussed in the sociology of religion and is reflected by incongruities among religious observers. Chaves (2010) illustrates this: “Ideas and practices exist as bits and pieces that come and go as situations change, producing many inconsistencies… Religious ideas, values, and practices generally are not congruent.”
@@ 2 Related Work
Recent NLP research has underscored the importance of narrative analysis in the study of human behavior and belief systems. While these studies often center around event and location schemas Wagner et al. (2025), character personalities, narrator intentions Piper et al. (2021); Zhu et al. (2023), or narrative structure Wagner et al. (2022), there remains a gap in tracking the evolution of characters, particularly their development. This gap is crucial, as understanding character evolution can provide deeper insight into how narratives convey complex human experiences.
@@
This study builds upon these foundations by analyzing character development, focusing on the protagonist’s described religious practices and their expressed religious belief systems.
@@
In the context of the Holocaust, Lassley (2015) examines religious transformation among Jewish survivors of the Holocaust, focusing on testimonies that describe a shift from being connected with religious traditions to abandoning belief in God completely. This work discusses how survivor testimonial information can serve as valuable sources regarding faith after the Holocaust.
@@ 3 Task Definition
Given a segmented testimony $(x_{1},x_{2},\ldots,x_{n})$, and a set of labels for practice and belief content $\mathcal{B}$ and $\mathcal{P}$, our framework outputs a sequence representing its religious trajectory $(\langle b_{1},p_{1}\rangle,\langle b_{2},p_{2}\rangle,\ldots,\langle b_{n},p_{n}\rangle)$, where $b_{j}\in\mathcal{B}\cup\{\texttt{None}\}$ and $p_{j}\in\mathcal{P}\cup\{\texttt{None}\}$. The None label is used for segments that do not contain belief or practice content.
@@ 4 Data and Annotation
We received 1000 Holocaust survivor testimony transcripts archived by the Shoah Foundation [^1] https://sfi.usc.edu/ (SF), originally recorded as interviews on video in English. The significance of this resource in the study and remembrance of the Holocaust cannot be overstated. As the number of surviving witnesses diminishes, there is an urgent need to find new approaches to engage with the extensive collection of testimonies, with computational methods offering interdisciplinary pathways to trace patterns of belief, resilience, identity, and trauma (Wagner et al., 2025; Tóth et al., 2022).
Due to the lack of a structured format, our first step is to segment the data using natural divisions created by question-answer pairs followed by merging segments with fewer than 10 words and dividing those exceeding 100 words. The method aims to create segments that contain just enough information to determine its relevance to the religious trajectory. This method yields a dataset with the majority of the segments being between 50-100 words and reasonable paragraph breaks, mostly keeping separate ideas distinct. However, some samples may contain multiple topics or conflicting religious signals. For example: “Saturday I would go, and I was encouraged to go to the synagogue. And Sunday, I was encouraged to go to church.”
Our analysis is grounded in comparing religious trajectories along their narrative timelines, which raises the question of how the chronological order of events and the narrative sequence align. Examining this relationship shows they are indeed closely aligned (see Appendix A).
@@ 4.1 Annotations
After segmenting the data, we randomly sample and annotate approximately 4000 segments. The first annotation assignment is for capturing descriptions of Jewish religious content. This is designed as a binary classification task – the annotators are directed to classify all segments that describe any Jewish religious practices or beliefs of the protagonist or indicate their absence. Complete guidelines are provided in Appendix B.
@@
This complex annotation task involves knowledge of history and the Jewish religion. Many segments can be ambiguous, rely on context, or require an understanding of the context. Therefore, there may be a lack of consensus among the annotators. All these factors contribute to the complexity of the task, which requires careful attention.
The annotation process was carried out by three annotators. We keep an overlap of 833 samples annotated by at least two annotators, for measuring inter-annotator agreement (IAA). To evaluate the IAA, we use Krippendorff’s alpha Krippendorff (2011) on the pairwise annotations, obtaining an average score of 0.74, indicating a substantial level of agreement. Appendix D presents the label distribution of the overlap set of annotations.
In case of disagreement within the overlap: if only two annotators reviewed the sample it was discarded from the dataset. Otherwise, the majority label was assigned. Due to the challenge of collecting annotations, the rest of the samples were annotated by a single annotator. Additionally, throughout the annotation process, we occasionally review the annotations and ask the annotators to revise their work when necessary.
@@
The second annotation task is multi-label and consists of identifying the class of each sample and determining the speaker’s valence toward it. Each segment belonging to the trajectory is independently categorized by its content – practice, belief, or both (see §1). The segments are further categorized by their valence toward the specific practice or belief they describe. Specifically, we map each sample to one or more of the classes or mark it as falsely recognized religious content.
@@
A segment describing a practice can be annotated with one of the classes – Active, Inactive or Other. The first two classes identify descriptions of participating in or abstaining from religious practices while Other represents descriptions that do not meet the criteria of Active or Inactive, or match both of the classes simultaneously. For belief, the classification schema is similar, with the classes being Positive, Negative, and Other. An example of the annotation process is illustrated in Figure 2, see Appendix C for the full guidelines.
@@
To construct the dataset for this task, we first run the religious content classifier to identify segments of interest. Since the majority of segments have no religious content, this step serves as initial filtering before the detailed annotation. Additionally, we use GPT-4o [^2] https://platform.openai.com/docs/models/gpt-4o version gpt-4o-2024-05-13 to identify samples that belong to the minority classes, specifically the Belief classes and Other-Practice. We then incorporate into the dataset all samples that were predicted to be in these minority classes. This approach helps balance the dataset by increasing the number of examples in the underrepresented classes, ensuring both positive and negative instances are well-represented. The resulting dataset is the one provided to the annotators.
@@
We collect 1,165 annotations and hold an overlap of 540 samples. The average IAA score for the second task is 0.48, indicating moderate agreement and suggesting that this classification task is challenging even for human annotators. We divide the data into three splits with equal proportions of each class. The test split is completely sampled from the overlap, to ensure the reliability of our results. Appendix D summarizes the label distribution for both annotation tasks.
@@ 5.1 Experimental setup
Extracting the religious trajectory from a segmented testimony consists of fine-tuning and prompting LLMs. The main steps are: (1) filter religious content, (2) identify practices and beliefs, and (3) measure the valence of the protagonist’s expressed attitude toward religion.
@@
The annotation process yields 674 positive samples with religious content and the rest negative. We fine-tune a RoBERTa-large Liu et al. (2019) classifier on this task, using a balanced dataset (1,348 samples), divided into three splits: train, validation, and test, by a ratio of 0.8, 0.1 and 0.1, respectively. [^3] Hyperparameters: 4 epochs, batch size=4, learning rate=1e-5, seed=5 To enhance the reliability of the model and reduce individuals’ bias in the evaluation, the entire test split was randomly sampled from the overlap data.
Evaluation of the model on the test set yields accuracy and recall scores of 97% and 94% respectively. Additionally, cross-validation over 10 random splits produces an average accuracy and recall of 91%.
Given the goal of this assignment to identify segments that form a node in the trajectory – i.e., contain relevant practice and belief descriptions, these results demonstrate that the RoBERTa classifier is a reliable tool for this purpose.
For modeling the trajectories, we prompt a series of GPT-4.1 models [^4] https://platform.openai.com/docs/models/gpt-4.1, version: 2025-04-14 on multiple settings and compare their performance to fine-tuned LoRA adapters for Mistral-7B-instruct, using data from the second annotation task. Appendix F contains the prompts we experiment with. Model selection is based on the macro average F1 score on the test set. For training, we use Mistralai’s mistral-finetune repository [^5] https://github.com/mistralai/mistral-finetune with their default hyperparameters and test multiple seeds [^6] Hyperparameters: Lora Rank=64, sequence length=16,384, batch size=1, learning rate=1e-4 number of micro-batches=1. Number of training steps=100, weight decay=0.1, pct-start=0.05. Seed for belief=1; for practice=5. The best performing model across all practice and belief classes is GPT4.1 with an average F1 of 0.55 for practice and 0.64 for belief. The fine-tuned Mistral-7B scores 0.43 and 0.45 for practice and belief, respectively. In addition, we experiment with GPT-4.1-mini, which results in 0.48 (practice) and 0.62 (belief). See Appendix G for the complete details on the multiple settings. Owing to limitations with large-scale API access, we use Mistral-7B to generate the full dataset of trajectories. The prompts we run follow the format of Self-Consistency Wang et al. (2022) and are carefully designed with guidance from Anthropic’s prompt generator. [^7] https://console.anthropic.com/login Considering the complicated annotation process and IAA, along with the models’ bias toward stereotypes in the context of religion Plaza-del Arco et al. (2024), these moderate F1 scores are unsurprising.
@@ 5.2 Evaluation
As we are not aware of any directly comparable reference for the trajectories. We therefore evaluate them against two reliable proxies, albeit not directly comparable ones: topic-modeling-based and thesaurus-based references.
@@
Ifergan et al. (2024) applied topic modeling using BERTopic Grootendorst (2022) to the same set of testimony transcripts. Their results include multiple topics that correspond to religious practices and beliefs. Using this prior knowledge, we define a mapping of the topics to the practice/belief labels, and treat the result as reference trajectories. The segments that were used for topic modeling are about three times longer than the segments we use, leading to relatively sparse sequences. The induced topics that we compare against are provided in Appendix H.
@@
The testimony transcripts from the SF were divided into segments of one minute each, indexed with a subset of keywords from the SF’s detailed thesaurus of  8,000 terms.888sfi.usc.edu/content/keyword-thesaurus We leverage the terms related to religious practices and beliefs, and map them to practice/belief labels to extract reference trajectories. The full list of addressed terms can be found in Appendix I.
@@
Both references have several limitations. First, the segments in the reference trajectories are larger than ours, leading them to contain information irrelevant to the practice and belief labels. Second, the reference sequences are sparse compared to the predicted ones. For the topics, our analysis of the topic model selects a single topic for each segment, rather than all the top relevant topics. For the thesaurus, despite the large number of terms, many segments lack labels for their matching terms. This may be a result of the exceedingly large set of keywords, which may have resulted in recall issues in the SF annotation. In addition, the label set of the references is partial, meaning that the majority of the topics and terms we address do not have a specific valence. Considering these points, the false positive rate of the predictions is unknown, making the references suitable only for measuring recall. An example illustrating the alignment of the references and predictions is shown in Figure 3.
@@
If $T$ is empty, return the number of segments in $R$, and if $R$ is empty, return 0.
We compare the sum of the distances to baseline trajectories. These baselines are artificial sequences designed by using our prior knowledge about the distribution (Figure 5) of religious content throughout the dataset; each is the same length as the predicted sequence. The different baselines for a given class are defined in Table 1.